OpenStudy (anonymous):

I don't understand the last question I asked. Someone help me? I don't know the square roots.

OpenStudy (anonymous):

which one post it,please.

OpenStudy (anonymous):

just remember quadratic formula if u find hard factoring

OpenStudy (anonymous):

solutions to 4x^2-12x+9=5?

OpenStudy (anonymous):

just use quadratic eq

OpenStudy (anonymous):

4(x^2 - 3x + 1)

OpenStudy (anonymous):

it'll have irrational roots

OpenStudy (anonymous):

Yes I know they are

OpenStudy (anonymous):

4x^2-12x+4=0 -b+-\[\sqrt{b^2-4ac}/2a\]

OpenStudy (amistre64):

4x^2 -12x +9 = 5 ; -9 4x^2 -12x = -4 ; /4 x^2 -3x = -1 ; +(3/2)^2 x^2 -3x +(3/2)^2 = -1 +(3/2)^2 ; compact the left side (x-3/2)^2 = -1 +(9/4) ; sqrt the sides x - (3/2) = +- sqrt(5/4) ; +(3/2) x = (3/2) +- sqrt(5)/2 \[x = \frac{3 \pm \sqrt{5}}{2}\]

OpenStudy (anonymous):

THANK YOU SO VERY MUCH!

OpenStudy (amistre64):

youre welcome :) there are a few ways to get to the same answer; but I like to spice it up every now and again...